One of two supplementary angles is 4 times as large as the other. What is the complement of the smaller angle?
Answer by Paul(988)

The supplement of an angle is 180-x
THe complement of an angle is 90-x
Equation:
4(90-x)=180-x
-3x=-180
x=60
Hence, the complement of the smaller angle is 90-60=30 degrees.
Paul.